# Policy Search, Retrieval, and Composition via Task Similarity in Collaborative Agentic Systems

### v2

@@ Abstract
Agentic AI aims to create systems that set their own goals, adapt proactively to change, and refine behavior through continuous experience. Recent advances suggest that, when facing multiple and unforeseen tasks, agents could benefit from sharing machine-learned knowledge and reuse policies that have already been fully or partially learned by other agents. However, how to query, select, and retrieve policies from a pool of agents, and how to integrate such policies remains a largely unexplored area. This study explores how an agent decides what knowledge to select, from whom, and when and how to integrate it in its own policy in order to accelerate its own learning. The proposed algorithm, Modular Sharing and Composition in Collective Learning (MOSAIC), improves learning in agentic collectives by combining (1) knowledge selection using performance signals and cosine similarity on Wasserstein task embeddings, (2) modular and transferable neural representations via masks, and (3) policy integration, composition and fine-tuning. MOSAIC outperforms isolated learners and global sharing approaches in both learning speed and overall performance, and in some cases solves tasks that isolated agents cannot. The results also demonstrate that selective, goal-driven reuse leads to less susceptibility to task interference. We also observe the emergence of self-organization, where agents solving simpler tasks accelerate the learning of harder ones through shared knowledge.
@@ 1 Introduction
The ability to solve complex, open-ended problems is desirable objective as agentic AI systems become increasingly central in AI architectures. In real-world open-ended settings, these systems face challenges in generalizing and adapting to an ever evolving stream of data and tasks (De Lange et al. 2022; Parisi et al. 2019; Lesort et al. 2020). While new lifelong learning algorithms (Kudithipudi et al. 2022) are being introduced to provide continual evolution and adaptation, when agents learn in isolation, they can only benefit from their limited experiences, unlike humans who benefit from collaboration and sharing experiences and skills. Recent studies suggest that a shift toward more collaborative types of learning, where agents selectively share and reuse modular knowledge from other agents, is key to reduce redundant learning, accelerate adaptation, and improve robustness (Tarale, Rietman, and Siegelmann 2025; Soltoggio et al. 2024; DARPA/MTO 2021).
@@
In the context of self-adaptation, reinforcement signals provide an effective tool to achieve continuous improvement in a wide range of domains (Sutton and Barto 2018; Jaderberg et al. 2017; Wang et al. 2025; Sun 2023). In addition, task similarities have been exploited to benefit learning in areas such as transfer learning and multi-task learning (Hendawy, Peters, and D’Eramo 2024; Wang et al. 2023; Sun et al. 2022; D’Eramo et al. 2020; Zeng et al. 2021; Espeholt et al. 2018). Lifelong learning approaches also leverage task similarity when previously learned tasks benefit the learning of new tasks, with an advantage often named forward transfer (Ben-Iwhiwhu et al. 2023; Kirkpatrick et al. 2017; Wolczyk et al. 2022; Kessler et al. 2022; Cannon and Şimşek 2024). However, many approaches to share knowledge, e.g., distributed RL (Sartoretti et al. 2019) or Federated Learning (Wen et al. 2023), often assume some form of centralization and uniformity of tasks, which is ill-suited to the requirements of agentic AI where each agents acts independently, asynchronously, and likely on a large variety of different tasks. Recent studies have begun to investigate sharing and reuse of policies in less constrained, evolving multi-agent contexts (Nath et al. 2023; Tarale, Rietman, and Siegelmann 2025; Rostami et al. 2017; Amani et al. 2023; Gao et al. 2021; Gerstgrasser, Danino, and Keren 2023). These developments reflect a growing recognition that scalable AI systems increasingly depend on selective, self-directed collaboration among autonomous learners.
@@
Knowledge and skill sharing among a set of independent and autonomous learners, while conceptually appealing, presents significant challenges. In particular, open research questions reflect the problems of how to identify what knowledge to acquire, from whom to acquire it, how such knowledge is best represented and and how to integrate it to manifest robust lifelong learning properties (DARPA/MTO 2021; Soltoggio et al. 2024). Furthermore, it is still unclear to what extent collaborative learning offers an advantage over isolated or centralized learning.
@@
To address these questions, this study introduces Modular Sharing and Composition in Collective Learning (MOSAIC) [^1] An implementation of MOSAIC is available at https://github.com/DMIU-ShELL/MOSAIC., a novel algorithm that proposes a solution to select appropriate knowledge across a collective of agents with the aim of implementing policy composition and reuse. The problem analysis and following ablation studies indicate that essential components of such a system are (1) knowledge selection via task similarity, (2) modular and transferable neural representations, and (3) policy integration, composition and fine-tuning.
@@
Specific algorithmic choices in MOSAIC to implement such components are Wasserstein embeddings for task similarity, transferable binary network masks for sharing, and learnable linear combinations of those to achieve integration and fine tuning. Alternative but equivalent methods could also be used to adapt MOSAIC to a variety of domains, e.g., LoRA-based modular composition for foundation models and LLMs (Hu et al. 2022; Huang et al. 2024; Zhong et al. 2024; Asadi et al. 2024; Prabhakar et al. 2025), and embedding methods or descriptors to express task alignment (Duan et al. 2017; Achille et al. 2019; Dou, Luo, and Yang 2022; Grover et al. 2018; Rakelly et al. 2019).
@@
Simulations indicate that selectively acquiring relevant knowledge is essential to effectively reuse policies. The composition of policies is best achieved with a similarity-driven and reward-aware weighting. The results show how communicating MOSAIC agents learn faster than isolated learners by reusing knowledge from peers on similar tasks. In some cases, communicating agents are able to solve tasks that they could not learn alone. In addition, the analysis reveals an implicit self-organization of learning into task curricula, where agents completing simpler tasks accelerate the progress of those tackling harder ones.
@@
In summary, MOSAIC supports the hypothesis that task-relevant policy transfer and reuse, guided by agent-centered optimization objectives, is feasible and effective with advantages over learning in isolation or sharing global parameters, which could lead to task interference. To the best of our knowledge, this is the first study that combines modular and transferable task-specific knowledge with similarity-based selection and integration. Such a combination of components are sufficient to significantly improve sample efficiency and learning through collective knowledge sharing.
@@ 2 Related Work
Recent work in modular knowledge reuse for multi-agent RL explores various approaches to enable efficient knowledge transfer. Federated learning methods (Yoon et al. 2021; Seo, Han, and Yoo 2025) communicate task-specific masks or binary representations. These approaches lack principled task similarity metrics for selective reuse. MOSAIC addresses this gap through Wasserstein-based similarity measures that guide knowledge selection based on performance signals.
Several frameworks enable parameter-efficient knowledge sharing through masking (Paul et al. 2025; Gao et al. 2021; Ge et al. 2023; Nath et al. 2023; Rostami et al. 2017; Gerstgrasser, Danino, and Keren 2023), prototype aggregation (Tan et al. 2022), or decentralized training (Douillard et al. 2024; Jaghouar, Ong, and Hagemann 2024). Communication-efficient methods reduce information exchange through knowledge preservation and selective sharing (Diskin et al. 2021; McMahan et al. 2017; Meng et al. 2025). These approaches rely on predefined sharing protocols or require synchronous updates. MOSAIC’s binary mask representation enables asynchronous, selective reuse and maintains modularity while minimizing communication overhead.
Policy composition methods like PaCo (Sun et al. 2022) interpolate within shared subspaces. These methods lack mechanisms for similarity-driven integration. Peer-to-peer frameworks (Sengupta et al. 2024; Alvar and Bajić 2019; Kaimakamidis, Mademlis, and Pitas 2024; Tarale, Rietman, and Siegelmann 2025; Yu, Vincent, and Schwager 2022; Nadiger, Kumar, and Abdelhak 2019; Lozano-Cuadra et al. 2025) and model-based approaches (Jiang, Narayanan, and Li 2021; Du et al. 2025; Shi et al. 2022) enable decentralized coordination. These frameworks do not combine performance-based selection with modular representations. MOSAIC integrates Wasserstein-based task similarity with binary mask modularity and similarity-driven linear policy combinations. MOSAIC demonstrates improved sample efficiency and self-organization in collaborative multi-agent systems. Further discussion of related works is provided in Appendix A.
@@ 3 Methodology
The following section describes the components of the MOSAIC algorithm. A high level illustration is provided in Figure 1. Pseudocode can be found in the Appendix E.
@@ 3.1 Policy Representations
Isolating task-specific knowledge in compact representations (Alet, Lozano-Perez, and Kaelbling 2018) is a key to make policies easy to transfer across agents. Several parameter isolation mechanisms from lifelong learning can achieve this goal (von Oswald et al. 2020; Hu et al. 2022; Rusu et al. 2016; Mallya and Lazebnik 2018).
This study adopts neural network masks (Ben-Iwhiwhu et al. 2023; Wortsman et al. 2020) that have shown strong results in lifelong reinforcement learning and supervised learning settings, and supports composition through linear combinations. This design enables integration with standard methods such as PPO (Schulman et al. 2017), DQN (Mnih et al. 2013), or A3C (Mnih et al. 2016).
Figure 1: High-level illustration of the main MOSAIC algorithmic steps. (A) A Wasserstein task embedding $v_{\tau}$ is computed from a SAR batch by Agent A0 (representing any agent in the collective). (B) Periodically, the agent A0 broadcasts a task embedding query (TEQ) to all known peers. (C) Peers send back a query response (QR) that contains their task’s $v_{\tau},r$, and the corresponding mask ID; (D) Agent A0 selects relevant embeddings using cosine similarity on the Wasserstein embeddings. (E) A0 further selects relevant embeddings using Criterion 1 and Criterion 2. (F) A0 sends mark requests (MR). (G) The contacted agents respond by sending the requested mask through a mask transfer (MTR). (H) Incoming masks from A2 and A3 are incorporated into A0. The training of the agent’s policy occurs in parallel (not represented in the figure).
Mask representations rely on a frozen backbone network $\Phi$ that is homogeneous across all agents. Each policy $\pi_{\tau}$ is implicitly parameterized through a sparse mask $\phi_{\tau}$, such that $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau})}$, where $\odot$ denotes the Hadamard product, and $g(\cdot)$ is a binarization function applied during the forward pass. The mask $\phi_{\tau}$ consists of a set of real-valued score vectors, one per layer of the shared backbone network $\Phi$, which select a sparse, task-specific subnetwork from $\Phi$ without modifying its parameters. Scores are binarized during the forward pass to select a discrete subnetwork from the backbone. The binarization is given by the element-wise thresholding function,
@@
where $g(\phi_{\tau})$ denotes the full binary mask and $\ell$ indexes the layers of the backbone network $\Phi$. During the backward pass, gradients are computed with respect to the real-valued mask parameters $\phi_{\tau}$, and updates are performed using the straight-through estimator (STE) (Wortsman et al. 2020) to enable learning through the non-differentiable binarization step.
@@ 3.2 Wasserstein Task Embeddings for Online Reinforcement Learning
To identify suitable policies for a given tasks, MOSAIC extends the Wasserstein Task Embedding (WTE) framework of Liu et al. (2025) to RL by computing embeddings over batches of state-action-reward (SAR) data collected online (see Figure 1(A)). For task $\tau$, the agent maintains an empirical task distribution
@@
where $x_{t}=(s_{t},a_{t},r_{t})$ and $x_{t}\sim\mathcal{D}_{\tau}$. Here, $\mathcal{D}_{\tau}$ denotes the replay buffer for task $\tau$, and each $x_{t}\in\mathbb{R}^{d}$ is a state-action-reward (SAR) tuple. The tuple dimensionality is $d=d_{s}+d_{a}+d_{r}$, where $d_{s}$, $d_{a}$, and $d_{r}$ are the dimensions of the state, action, and scalar reward, respectively. We use $N=128$ samples in our experiments. $\delta_{x_{t}}$ denotes the Dirac measure centered at the sample $x_{t}$. A fixed synthetic reference distribution $\mu_{0}$ is defined as,
@@
This mapping places each task $\tau$ in a shared latent space, where metrics can be used to quantify task relationships.
@@ 3.3 Knowledge Selection and Sharing
Agents communicate maintaining a list of IP addresses and ports for other agents, which allows them to be located anywhere on the Internet. Each agent $i\in\mathcal{I}$ can directly communicate with all other agents $j\in\mathcal{I},\ j\neq i$ at any time.
@@
At any time, an agent $i$ can initiate a task embedding query (TEQ) by broadcasting its current task embedding $v_{i}$, iteration performance $r_{i}$, and its IP/port (see Figure 1(B)(C)). Upon receiving the TEQ, each peer agent $j\neq i$ responds with its most recent task embedding $v_{j}$ and associated performance score $r_{j}$.
@@
$d_{\text{cos}}(v_{i},v_{j})=\frac{\langle\text{vec}(v_{i}),\text{vec}(v_{j})\rangle}{|\text{vec}(v_{i})|\cdot|\text{vec}(v_{j})|},\quad d_{\text{cos}}\in[-1,1].$ (6)
@@
$\mathbb{I}_{\text{align}}(i,j)$ $=\begin{cases}1&\text{if }d_{\text{cos}}(v_{i},v_{j})>\theta,\\ 0&\text{otherwise}\end{cases}$ (Criterion 1)
$\mathbb{I}_{\text{perf}}(i,j)$ $=\begin{cases}1&\text{if }r_{j}>r_{i},\\ 0&\text{otherwise}\end{cases}$ (Criterion 2)
with $\theta=0.5$ in all experiments. The first criterion promotes semantic alignment between tasks, revealing latent similarities that may lead to effective transfer (see Section 4.4). The second ensures that agents only acquire masks from peers whose performance exceeds their own, avoiding noisy or under-trained policies that can degrade performance. Criterion 2 guards against the frequent false positives seen when comparing poorly trained policies that yield deceptively similar embeddings.
The two phases and the selection criteria also result in a bandwidth-efficient approach as policies are transferred only when considered potentially useful. The peer masks that pass both criteria are stored as the set $P_{c+1}=\{\phi_{1},...,\phi_{K}\}$, where each $\phi_{k}$ denotes a mask received from a selected peer.

### v3

@@ Abstract
Agentic AI aims to create systems that set their own goals, adapt proactively to change, and refine behavior through continuous experience. Recent advances suggest that, when facing multiple and unforeseen tasks, agents could benefit from sharing machine-learned knowledge and reusing policies that have already been fully or partially learned by other agents. However, how to query, select, and retrieve policies from a pool of agents, and how to integrate such policies remains a largely unexplored area. This study explores how an agent decides what knowledge to select, from whom, and when and how to integrate it in its own policy in order to accelerate its own learning. The proposed algorithm, Modular Sharing and Composition in Collective Learning (MOSAIC), improves learning in agentic collectives by combining (1) knowledge selection using performance signals and cosine similarity on Wasserstein task embeddings, (2) modular and transferable neural representations via masks, and (3) policy integration, composition and fine-tuning. MOSAIC outperforms isolated learners and global sharing approaches in both learning speed and overall performance, and in some cases solves tasks that isolated agents cannot. The results also demonstrate that selective, goal-driven reuse leads to less susceptibility to task interference. We also observe the emergence of self-organization, where agents solving simpler tasks accelerate the learning of harder ones through shared knowledge.
Codehttps://github.com/DMIU-ShELL/MOSAIC
@@ 1 Introduction
The ability to solve complex, open-ended problems is an increasingly desirable objective as agentic AI principles become central in AI architectures. In open-ended real-world settings, these systems face challenges in generalizing and adapting to an ever-evolving stream of data and tasks (De Lange et al. 2022; Parisi et al. 2019). Although new lifelong learning algorithms (Kudithipudi et al. 2022) are being introduced to provide continual evolution and adaptation, when agents learn in isolation, they can only benefit from their limited experiences, unlike humans who benefit from collaboration and sharing of experiences and skills. Recent studies suggest that a shift towards more collaborative types of learning, where agents selectively share and reuse modular knowledge from other agents, is key to reducing redundant learning, accelerating adaptation, and improving robustness (Tarale, Rietman, and Siegelmann 2025; Soltoggio et al. 2024; DARPA/MTO 2021).
@@
In the context of self-adaptation, reinforcement signals provide an effective tool to achieve continuous improvement in a wide range of domains (Sutton and Barto 2018; Jaderberg et al. 2017). In addition, task similarities have been exploited to benefit learning in areas such as transfer learning and multi-task RL (Hendawy, Peters, and D’Eramo 2024; Sun et al. 2022; D’Eramo et al. 2020). Lifelong learning approaches also leverage task similarity when previously learned tasks benefit the learning of subsequent tasks, with an advantage often named forward transfer (Ben-Iwhiwhu et al. 2023; Kirkpatrick et al. 2017). However, many approaches to sharing knowledge, e.g., distributed RL (Sartoretti et al. 2019) or Federated Learning (Yoon et al. 2021), often assume some form of centralization and uniformity of tasks, which is ill-suited to the requirements of agentic AI where each agent acts independently, asynchronously, and likely on a large variety of different tasks. Recent studies have begun to investigate the sharing and reuse of policies in evolving multi-agent contexts (Nath et al. 2023; Tarale, Rietman, and Siegelmann 2025; Gerstgrasser, Danino, and Keren 2023). These developments reflect a growing recognition that scalable AI systems increasingly depend on self-directed selective collaboration among autonomous learners.
@@
Knowledge and skill sharing among a set of independent and autonomous learners, while conceptually appealing, presents significant challenges. In particular, open research questions reflect the problems of how to identify what knowledge to acquire, from whom to acquire it, how such knowledge is best represented, and how to integrate it to manifest robust lifelong learning properties (DARPA/MTO 2021; Soltoggio et al. 2024). Furthermore, it is still unclear to what extent collaborative learning offers an advantage over isolated or centralized learning.
@@
To address these questions, this study introduces Modular Sharing and Composition in Collective Learning (MOSAIC), which describes how agents can select appropriate knowledge from peers with the aim of implementing policy composition and reuse. Problem analysis and the following ablation studies indicate that essential components of such a system are (1) knowledge selection through task similarity, (2) modular and transferable neural representations, and (3) policy integration, composition, and fine-tuning.
@@
Specific algorithmic choices in MOSAIC to implement such components are Wasserstein embeddings for task similarity, transferable binary network masks for sharing, and learnable linear combinations of those to achieve integration and fine tuning. In principle, analogous methods could be used to adapt MOSAIC to a variety of domains, e.g., LoRA-based modular composition for foundation models and LLMs (Hu et al. 2022; He et al. 2022) and embedding methods or descriptors to express task alignment (Achille et al. 2019; Grover et al. 2018; Rakelly et al. 2019).
@@
Simulations indicate that selectively acquiring relevant knowledge is essential to effectively reuse policies. The composition of policies is best achieved with a similarity-driven and reward-aware weighting. The results show how communicating MOSAIC agents learn faster than isolated learners by reusing knowledge from peers on similar tasks. In some cases, communicating agents can solve tasks that they could not learn alone. In addition, the analysis reveals an implicit self-organization, where agents discover and exploit curriculum structures in the available knowledge, resulting in agents building on skills hierarchically, from simpler tasks to learn more complex ones.
@@
In summary, MOSAIC supports the hypothesis that task-relevant policy transfer and reuse, guided by agent-centered optimization objectives, is feasible and effective with advantages over learning in isolation or sharing global parameters, which could lead to task interference. To the best of our knowledge, this is the first study that combines modular and transferable task-specific knowledge with similarity-based selection and integration. Such a combination of components is sufficient to significantly improve sample efficiency and learning through collective knowledge sharing.
@@ 2 Related Work
Recent work in modular knowledge reuse for multi-agent RL explores various approaches to enable efficient knowledge transfer. Federated learning methods (Yoon et al. 2021) communicate task-specific masks or binary representations. These approaches lack principled task similarity metrics for selective reuse. MOSAIC addresses this gap through Wasserstein-based similarity measures that guide knowledge selection based on performance signals.
Several frameworks allow parameter-efficient knowledge sharing through masking (Ge et al. 2023; Nath et al. 2023; Gerstgrasser, Danino, and Keren 2023), prototype aggregation (Tan et al. 2022), or decentralized training (Douillard et al. 2024; Jaghouar and Hagemann 2024). Communication-efficient methods reduce information exchange through knowledge preservation and selective sharing (McMahan et al. 2017). These approaches rely on predefined sharing protocols or require synchronous updates. MOSAIC’s binary mask representation enables asynchronous, selective reuse and maintains modularity while minimizing communication overhead.
Policy composition methods like PaCo (Sun et al. 2022) interpolate within shared subspaces. These methods lack mechanisms for similarity-driven integration. Peer-to-peer frameworks (Tarale, Rietman, and Siegelmann 2025; Yu, Vincent, and Schwager 2022) and model-based approaches (Jiang, Narayanan, and Li 2021) enable decentralized coordination. These frameworks do not combine selection with modular representations. MOSAIC combines Wasserstein-based task similarity with binary mask modularity and similarity-driven linear policy combinations. Further discussion of related works is provided in Appendix B.
@@ 3 Methodology
The following section describes the components of the MOSAIC algorithm. A high level illustration is provided in Figure 1. Pseudocode can be found in the Appendix F.
@@ 3.1 Policy Representations
Isolating task-specific knowledge in compact representations (Alet, Lozano-Perez, and Kaelbling 2018) is a key to making policy transfer easier and more lightweight across agents. Lifelong learning parameter isolation approaches can achieve this goal (Rusu et al. 2016; Mallya and Lazebnik 2018; Wortsman et al. 2020; Ben-Iwhiwhu et al. 2023).
This study adopts neural network masks (Ben-Iwhiwhu et al. 2023; Wortsman et al. 2020) that have shown strong results in lifelong reinforcement learning and supervised learning settings, and supports composition through linear combinations of mask modules. PPO is used for the experiments in this paper; however, the proposed modular composition mechanism is, in principle, agnostic to the choice of RL algorithm.
Figure 1: High-level illustration of the main MOSAIC algorithmic steps. (A) A Wasserstein task embedding $v_{\tau}$ is computed from a SAR batch by Agent A0 (representing any agent in the collective). (B) Periodically, the agent A0 broadcasts a task embedding query (TEQ) to all known peers. (C) Peers send back a query response (QR) that contains their task’s $v_{\tau},r$, and the corresponding mask ID; (D) Agent A0 selects relevant embeddings using cosine similarity on the Wasserstein embeddings. (E) A0 further selects relevant embeddings using Criterion 1 and Criterion 2. (F) A0 sends mask requests (MR). (G) The contacted agents respond by sending the requested mask through a mask transfer (MTR). (H) Incoming masks from A2 and A3 are incorporated into A0. The training of the agent’s policy occurs in parallel (not represented in the figure).
Mask representations rely on a frozen backbone network $\Phi$ that is homogeneous across all agents. Each policy $\pi_{\tau}$ is implicitly parameterized through a sparse mask $\phi_{\tau}$, such that $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau})}$, where $\odot$ denotes the Hadamard product, and $g(\cdot)$ is a binarization function applied during the forward pass. The mask $\phi_{\tau}$ consists of a set of real-valued score vectors, one per layer of the shared backbone network $\Phi$, which select a sparse task-specific subnetwork from $\Phi$ without modifying its parameters. Scores are binarized during the forward pass to select a discrete subnetwork from the backbone, given by the thresholding function,
@@
where $g(\phi_{\tau})$ denotes the complete binary mask and $\ell$ indexes the layers of the backbone network $\Phi$. During the backward pass, gradients are computed with respect to the real-valued mask parameters $\phi_{\tau}$, and updates are performed using the straight-through estimator (STE) (Wortsman et al. 2020) to enable learning through the non-differentiable binarization step.
@@ 3.2 Wasserstein Task Embeddings for Online Reinforcement Learning
To identify suitable policies for a given task, MOSAIC extends Liu et al. (2025) to RL by computing embeddings over batches of state-action-reward (SAR) data collected online (see Figure 1(A)). For task $\tau$, the agent maintains an empirical task distribution
@@
where $x_{t}=(s_{t},a_{t},r_{t})$ and $x_{t}\sim\mathcal{D}_{\tau}$. Here, $\mathcal{D}_{\tau}$ denotes the replay buffer for $\tau$, and each $x_{t}\in\mathbb{R}^{d}$ is a state-action-reward (SAR) tuple. The tuple dimensionality is $d=d_{s}+d_{a}+d_{r}$, where $d_{s}$, $d_{a}$, and $d_{r}$ are the dimensions of the state, action, and scalar reward, respectively. We use $N=128$ samples in our experiments. $\delta_{x_{t}}$ denotes the Dirac measure centered on the sample $x_{t}$. A fixed synthetic reference distribution $\mu_{0}$ is defined as
@@
This mapping places each task $\tau$ in a shared latent space, where metrics can be used to quantify task relationships. A moving average is maintained to smooth fluctuations as the agent’s policy evolves, $v_{t+1}^{\tau}=\frac{1}{2}(v_{t}^{\tau}+v_{t+1}^{\tau})$.
@@ 3.3 Knowledge Selection and Sharing
Agents communicate by maintaining a list of IP addresses and ports for other agents, which allows them to be located anywhere on the Internet. Each agent $i\in\mathcal{I}$ can communicate directly with all other agents $j\in\mathcal{I},\ j\neq i$ at any time.
@@
At any time, an agent $i$ can initiate a task embedding query (TEQ) by broadcasting its current task embedding $v_{i}$, iteration performance $\overline{r}_{i}$, and its IP/port (see Figure 1(B)(C)). Upon receiving the TEQ, each peer agent $j\neq i$ responds with its most recent task embedding $v_{j}$ and associated performance $\overline{r}_{j}$.
@@
$\text{cos}(v_{i},v_{j})=\frac{\langle\text{vec}(v_{i}),\text{vec}(v_{j})\rangle}{|\text{vec}(v_{i})|\cdot|\text{vec}(v_{j})|},\quad\text{cos}\in[-1,1].$ (6)
@@
$\mathbb{I}_{\text{align}}(i,j)$ $=\begin{cases}1&\text{if }\text{cos}(v_{i},v_{j})>\theta,\\ 0&\text{otherwise}\end{cases}$ (Criterion 1)
$\mathbb{I}_{\text{perf}}(i,j)$ $=\begin{cases}1&\text{if }\overline{r}_{j}>\overline{r}_{i},\\ 0&\text{otherwise}\end{cases}$ (Criterion 2)
with $\theta=0.5$ in all experiments. The first criterion promotes semantic alignment between tasks, revealing latent similarities that can lead to effective transfer (see Section 4.4). The second ensures that agents only acquire masks from peers whose performance exceeds their own, avoiding noisy or under-trained policies that can degrade performance. Criterion 2 guards against the frequent false positives seen when comparing poorly trained policies that yield deceptively similar embeddings.
The two phases and the selection criteria also result in a bandwidth-efficient approach, since policies are transferred only when considered potentially useful. The peer masks that pass both criteria are stored as the set $P_{c+1}=\{\phi_{1},...,\phi_{K}\}$, where each $\phi_{k}$ denotes a mask received from a selected peer.